OpenStudy (anonymous):

A substance, A2B, has the composition by mass of 60% A and 40% B. What is the composition of AB2 by mass?

OpenStudy (anonymous):

Draw a reaction profile for reaction 2N2+O2+heat→2N2O, must include proper labels for each axis, reactants, products, indication for activation Energy and transition state

OpenStudy (aaronq):

Use an arbitrary mass, 100 g is an easy number to work with. 60% of 100 g is 60 g, there are two A's. Each A is 30 g 40 g is B, and there is only one, so B is 40 g. \(AB_2\), would have a mass of 30 g + 2*40 g = 110 g The new percent by mass composition of A is: \(\dfrac{30g}{110g}*100\%=27.3\%\) The new percent by mass composition of B is: \(\dfrac{80g}{110g}*100\%=72.7\%\)